OpenStudy (anonymous):

Differentiate y = 7^x,

OpenStudy (anonymous):

7^x log 7

OpenStudy (anonymous):

u kno diff of a^x is a^x loga

OpenStudy (anonymous):

tyvm!

OpenStudy (anonymous):

\[\huge \color{green}{\frac{d}{dx}(a^x) = a^x \times Log(a)}\]